# Attention Mechanisms Don’t Learn Additive Models: Rethinking Feature Importance for Transformers

### v1

@@ E.1.2 Post-hoc fitting of surrogate models
For the results in Figure 4, we then sample $10000$ new samples from this dataset and forward them through the trained transformers. The model log-odds score together with the feature vectors are used to train the different surrogate models, linear model, GAM, and SLALOM. For the linear model, we fit an OLS on the log-odds returned by the model. We use the word counts for each of the 10 tokens as a feature vector. The GAM provides the possibility to assign each token a different weight according to its position in the sequence. To this end, we use a different feature vector of length $30\cdot 10$. Each feature corresponds to a token and a position and is set to one if the token $i$ is present at this position, and set to 00 otherwise. We then fit a linear model using regularized (LASSO) least squares with a small regularization constant of $\lambda=0.01$.
E.2 Recovering SLALOMs from observations
We resort to a second synthetic dataset to study the recovery property for the SLALOM. To find a realistic distribution of scores, we compute a BoW importance scores for input tokens of the BERT model on the IMDB dataset by counting the class-wise occurrence probabilities. We select $200$ tokens randomly from this dataset. We use these scores as value scores $v$ but multiply them by a factors of $2$ as many words have very small BoW importances. In realistic datasets, we observed that value scores $v$ are correlated with the importance scores $s$. Therefore, we sample
which results in the value/importance distribution given in Figure 7. We assign each word an equal occurrence probability and sample sequences of words at uniformly distributed lengths in random $[1,30]$. After a sequence is sampled, labels are subsequently sampled according to the log-odds ratio of the SLALOM. We train the transformer models on $20\cdot 20000$ samples. When using a smaller vocabulary size, we only sample the sequences out of the first $\lvert\mathcal{V}\rvert$ possible tokens.
@@ E.2 Recovering SLALOMs from observations
E.3 Training Details for Real-World data experiments
Training details. In these experiments, we use the IMDB [32] and Yelp [4] datasets to train transformer models on. Specifically, the results in Table 2(a) are obtained by training 2-layer versions of BERT, DistilBERT and GPT-2 with on $5000$ samples from the IMDB dataset for $1$ epoch, respectively. We did not observe significant variation in terms of number of layers, so we stick to the simpler models for the finial experiments. For the experiments in Table 2(b) we train $1$-layer versions of the above models for $1$ epoch on $25000$ samples of the Yelp dataset. We report the accuracies of these models in Table 4 and additional hyperparameters in Table 5.
@@ E.3 Training Details for Real-World data experiments
SLALOM vs. Human Attention. The Yelp Human Attention (HAT) [45] dataset consists of samples from the original Yelp review dataset, where for each review real human annotators have been asked to select the words they deem important for the underlying class (i.e. positive or negative). This results in binary attention vectors, where each word either is or is not important according to the annotators. Since each sample is processed by multiple annotators, we use the consensus attention map as defined in Sen et al. [45] to aggregate them into one attention vector per sample. Since HAT, unlike SLALOM, operates on a word level, we map each word’s human attention to each of its tokens (according to the employed model’s specific tokenizer).
@@
IMDB 0.84 0.74 0.81
Yelp 0.91 0.89 0.91
Table 4: Accuracies of models used in this paper. For IMDB ($|trainset|=5000$), we use 2-layer versions of the models. For Yelp, ($|trainset|=25000$), we use 1-layer versions of the models. For both datasets, the $|testset|=5000$. The models are trained for $1$ epoch after which we found the accuracy of the model on the test set to be converged.
@@
epochs 1
dataset size 5000
number of heads 6
@@ F.2 Fitting SLALOM on Transformers trained on data following the SLALOM distribution
We report Mean-Squared Errors in the logit-space and the parameter-space between original SLALOM scores and recovered scores. The logit output are evaluated on a test set of $200$ samples that are sampled from the original SLALOM. We provide these quantitative results in Table 7 for logit scores the result are very small Table 8. In logits the differences are negligibly small, and seem to decrease further with more layers. This finding highlight that a) transformers with more layers still easily fit SLALOMs and such model can be recovered in parameters space. The results on the MSE in parameter space show no clear trend, but are relatively small as well (with the largest value being MSE=$0.015$ (note that results in the table are multiplied by a factor of $100$ for readability). Together with our quantitative results in Figure 4, this highlight that SLALOM has effective recovery properties.
@@
Figure 8: SLALOM describes outputs of transformer models well. Fitting SLALOM to the outputs of the models shown in Figure 3. Results on additional models. Despite having $C/2{=}15\times$ more parameters than the SLALOM model, the GAM model does not describe the output as accurately. We provide quantitative results in Table 6.
F.3 Comparing SLALOM-SGD and SLALOM-Ana
We test our two fitting algorithms for different context lengths and vocabulary sizes. We report averages over 15 runs and present the results in Figure 9. We find that in the parameter space, SLALOM-Ana and SLALOM-SGD yield similar approximation quality. On the logit outputs, the SGD method results in a lower, almost constant approximation error whereas this error increases drastically for high context lengths when using SLALOM-Ana. This is not unexpected, as the analytical solution only uses short sequences to estimate the model. We therefore resort to SLALOM-SGD in our final study.
(a) Effect of vocabulary size $\lvert\mathcal{V}\rvert$
(b) SLALOM-Ana vs. SLALOM-SGD for different $C$
Figure 9: SLALOM-Ana vs. SLALOM-SGD for different context lengths $C$ and vocabulary sizes $\lvert\mathcal{V}\rvert$
F.4 Real-World Data
We obtain SLALOM explanations for real-world data using the procedure outlines in Algorithm 1 with sequences of length $n=3$.
In Figure 10 we show a sample from the Yelp HAT dataset. To visualize the attention of SLALOM, we train a $1$-layer BERT model with $12$ attention heads for $1$ epoch on reviews from the Yelp dataset. After fitting SLALOM on top of the resulting model, we can extract the importance scores given to each token in the sample. We can see that the SLALOM scores manage to identify many of the tokens which real human annotators also deemed important for this review to be classified as positive.
Figure 11 shows the full results from the sample used in Figure 5, where we only visualized a choice of words for readability purposes. For this, we train $2$-layer, $12$-head BERT and GPT-2 models on the IMDB dataset for $1$ epoch, respectively. After fitting SLALOM on these models, we use it to explain a movie review taken from the dataset, visualizing value scores $v$ against importance scores $s$.
(a) SLALOM importance scores
(b) Human Attention
Figure 10: SLALOM importance scores compared to human attention. Fitting SLALOM to the outputs of a 1-layer BERT model trained on Yelp reviews. Many words deemed important by human annotators are likewise highlighted by SLALOM.
@@ F.4 Real-World Data
F.5 Removal Benchmarks
@@
It is important to verify that SLALOM scores are competitive to other methods in classical explanation benchmarks as well. We therefore ran the classical removal and insertion benchmarks with SLALOM compared to baselines such as LIME, SHAP, Grad [48], and Integrated Gradients (IG, [51]). For the insertion benchmarks, the tokens with the highest attributions are inserted to quickly obtain a high prediction score to the target class. For the deletion benchmark, the tokens with the highest attributions are deleted from the sample to obtain a low score for the target class. We subsequently delete/insert more tokens and compute the “Area Over the Perturbation Curve” (AOPC) as in DeYoung et al. [16], which should be high for deletion and low for insertion. In addtion to the deletion results in Table 2(c), the insertion results are shown in Table 9. Our linear SLALOM scores perform par with LIME and SHAP in this benchmark. For surrogate techniques (LIME, SHAP, SLALOM) we use 5000 samples each. As a model, we use a 6-layer models for each of the named architectures.
LM values $v$ importances $s$ lin. SLALOM LIME SHAP IG Grad
GPT-2 0.088 $\pm$ 0.061 0.531 $\pm$ 0.209 0.088 $\pm$ 0.061 0.089 $\pm$ 0.062 0.100 $\pm$ 0.071 0.617 $\pm$ 0.191 0.559 $\pm$ 0.191
BERT 0.011 $\pm$ 0.005 0.293 $\pm$ 0.199 0.010 $\pm$ 0.004 0.009 $\pm$ 0.004 0.010 $\pm$ 0.006 0.482 $\pm$ 0.275 0.497 $\pm$ 0.280
DistilBERT 0.041 $\pm$ 0.020 0.442 $\pm$ 0.336 0.040 $\pm$ 0.019 0.041 $\pm$ 0.021 0.044 $\pm$ 0.025 0.439 $\pm$ 0.335 0.343 $\pm$ 0.319
@@
Table 9: Area-Over Perturbation Curve (insertion, lower is better)
F.6 Error Analysis for non-transformer models
@@
We also investigate the behavior of SLALOM for models that do not precisely follow the architecture described in the Analysis section of this paper. In the present work, we consider an attribution method that is specifically catered towards the transformer architecture, which is the most prevalent in sequence classification. Indeed, we observe mid-sized transformer models like BERT to be prevalent in sequence classification. On the huggingface hub, among the 10 most downloaded models, 9 are BERT-based and the remaining one is another transformer with around 33M parameters (MiniLM). We advise caution when using our model when the type of underlying LM is unknown. In this case, model-agnostic interpretability methods may be preferred. However, we investigate this issue further: We applied our SLALOM approach to a simple, non-transformer sequence classification model on the IMDB dataset, which is a three-layer feed-forward network based on a TF-IDF representation of the inputs. We compute the insertion and deletion Area-over-perturbation-curve metrics that are given in Table 10.
SHAP LIME lin. SLALOM
@@
These results show that due to its general expressivity, the SLALOM model also succeeds to provide explanations for non-transformer models that outperform LIME and SHAP in the removal and insertion tests. We also invite the reader to confer Table 13 and Section F.8, where we show that SLALOM can predict human attention for large models, including the non-transformer Mamba model [21].
F.7 Runtime analysis

### v2

@@ E.2 Post-hoc fitting of surrogate models
For the results in Figure 4(a,b), we then sample $10000$ new samples from the linear synthetic dataset (having the same distribution as the training samples) and forward them through the trained transformers. The model log odds score together with the feature vectors are used to train the different surrogate models, linear model, GAM, and SLALOM. For the linear model, we fit an OLS on the log odds returned by the model. We use the word counts for each of the 10 tokens as a feature vector. The GAM provides the possibility to assign each token a different weight according to its position in the sequence. To this end, we use a different feature vector of length $30\cdot 10$. Each feature corresponds to a token and a position and is set to one if the token $i$ is present at this position, and set to 00 otherwise. We then fit a linear model using regularized (LASSO) least squares with a small regularization constant of $\lambda=0.01$ because the system is numerically unstable in its unregularized form.
E.3 Recovering SLALOMs from observations
To obtain the results in Figure 4(c,d), we train the transformer models on $20\cdot 20000$ samples from the second synthetic dataset (SLALOM dataset). When using a smaller vocabulary size, we only sample the sequences out of the first $\lvert\mathcal{V}\rvert$ possible tokens (but keeping the ground truth scores identical).
@@ E.3 Recovering SLALOMs from observations
E.4 Training Details for real-world data experiments
Training details. In these experiments, we use the IMDB (Maas et al., 2011) and Yelp (Asghar, 2016) datasets to train transformer models on. Specifically, the results in Table 2(a) are obtained by training 2-layer versions of BERT, DistilBERT and GPT-2 with on $5000$ samples from the IMDB dataset for $5$ epochs, respectively. We did not observe significant variation in terms of number of layers, so we stick to the simpler models for the final experiments. For the experiments in Table 2(b) we train and use $6$-layer versions of the above models for $5$ epochs on $5000$ samples of the Yelp dataset. We report the accuracies of these models in Table 5 and additional hyperparameters in Table 6.
@@ E.4 Training Details for real-world data experiments
SLALOM vs. Human Attention. The Yelp Human Attention (HAT) (Sen et al., 2020) dataset consists of samples from the original Yelp review dataset, where for each review real human annotators have been asked to select the words they deem important for the underlying class (i.e. positive or negative). This results in binary attention vectors, where each word either is or is not important according to the annotators. Since each sample is processed by multiple annotators, we use the consensus attention map as defined in Sen et al. (2020), requiring agreement between annotators for a token to be considered important to aggregate them into one attention vector per sample. Since HAT, unlike SLALOM, operates on a word level, we map each word’s human attention to each of its tokens (according to the employed model’s specific tokenizer).
@@
IMDB 0.88 0.90 0.74
Yelp 0.86 0.88 0.88
Table 5: Accuracies of models used in this paper. For IMDB ($|trainset|=5000$), we use 2-layer versions of the models. For Yelp, ($|trainset|=5000$), we use 6-layer versions of the models. For both datasets, the $|testset|=100$. The models are trained for $5$ epochs after which we found the accuracy of the model on the test set to have converged.
@@
epochs 5
dataset size used 5000
number of heads 12
number of layers 2 (IMDB), 6 (Yelp)
@@ F.1 Fitting SLALOM as a Surrogate to Transformer outputs
We provide an additional emprical counterexample for why GAMs cannot describe the transformer output in Figure 9. The example provides additional intuition for why the GAM is insufficient to describe transformers acting like a weighted sum of token importances.
@@ F.2 Fitting SLALOM on Transformers trained on data following the SLALOM distribution
Input Sequence BERT score GAM weight assignments SLALOM weight assignments
[…]
worst perfect 0.0 ↯  Expected GAM output
[…]
Figure 9: A simple empirical counterexample for why GANs cannot describe transformer output. We report rounded scores by a real 4-layer BERT model (similar behavior was observed for other layers/architectures) and iteratively fit the GAM $F({\bm{t}})=\sum_{t_{i}\in{\bm{t}}}w_{i}(t_{i})$ to match observed outputs on the two tokens “perfect” and “worst”. We quickly arrive at a contradiction for the GAM. On the contrary, we can assign SLALOM scores that model this behavior with minor error. Because transformers behave like a weighted sum of importances, GAMs are insufficient to model their behavior. In conjunction with Figure 4(a,b) this underlines that GAMs and linear models are insufficient as surrogates.
We report Mean-Squared Errors in the logit-space and the parameter-space between original SLALOM scores and recovered scores. The logit output are evaluated on a test set of $200$ samples that are sampled from the original SLALOM. We provide these quantitative results in Table 8 in Table 9 for the parameter space. In logits the differences are negligibly small, and seem to decrease further with more layers. This finding highlight that a) transformers with more layers still easily fit SLALOMs and such model can be recovered in parameters space. The results on the MSE in parameter space show no clear trend, but are relatively small as well (with the largest value being MSE=$0.015$ (note that results in the table are multiplied by a factor of $100$ for readability). Together with our quantitative results in Figure 4(c,d), this highlights that SLALOM has effective recovery properties.
@@
Figure 10: SLALOM describes outputs of transformer models well. Fitting SLALOM to the outputs of the models shown in Figure 3 using the synthetic dataset. We show results for a sequence containing repetitions of the token “perfect”. Note however that the models were trained on a larger dataset of random sequences samples as described in Section E.1.1, but these sequences were chosen for visualization purposes. Results on additional models. Despite having $C/2{=}15\times$ more parameters than the SLALOM model, the GAM model does not describe the output as accurately. We provide quantitative results in Table 7.
F.3 Additional Results on Real-World Data
We obtain SLALOM explanations for real-world data using the procedure outlined in Algorithm 1 (SLALOM-eff) with sequences of length $n=2$ and with Algorithm 2 (SLALOM-fidel) removing up to $5$ tokens that we compare with Naive-Bayes scores and Human Attention.
F.3.1 Additional Qualitative Results
Figure 12 shows the full results from the sample used in Figure 5, where we only visualized a choice of words for readability purposes. After running SLALOM-eff on our trained IMDB models, we use to explain a movie review taken from the dataset, visualizing value scores $v$ against importance scores $s$.
F.3.2 Correlation with Naive-Bayes Scores
We compare the scores obtained with SLALOM with the scores obtained with other methods in Table 10, obtaining scores that are reliable with SLALOM-eff (value scores and linear scores) in particular. While SHAP achieves higher correlation on BERT, SLALOM achieves higher correlation than LIME and SHAP on all models and higher correlations than LRP for GPT-2 while obtaining slightly inferior values for the BERT-based architectures.
F.3.3 Human Attention
In Figure 11, we show qualitative results for a sample from the Yelp-HAT dataset. After fitting SLALOM on top of the resulting model, we can extract the importance scores given to each token in the sample. We can see that the SLALOM scores manage to identify many of the tokens which real human annotators also deemed important for this review to be classified as positive. We also show qualitative results for the other methods. However, we suggest caution when interpreting explanations visually without ground truth. We argue that (1) theoretical properties of explanations (2) comparing to a known ground truth as well as (3) consideration of metrics from different domains, e.g., faithfulness, human perspective, are required to allow for a comprehensive evaluation. This is the approach taken in our work.
We show a quantitative comparison of the scores obtained with SLALOM with the scores obtained with other methods on the comparison with Human-Attention in Table 11.
(a) Human annotation
(b) SLALOM-eff
(c) SHAP
(d) LIME
(e) AttnLRP
Figure 11: Qualitative comparisons of attribution maps. We provide attribution maps for the different techniques in this figure. Many words deemed important by human annotators are likewise highlighted by SLALOM and other techniques.
@@ F.3.3 Human Attention
SLALOM-fidel SLALOM-eff
LM values $v$ importances $s$ lin. SLALOM values $v$ importances $s$ lin. SLALOM
DistilBERT 0.602 $\pm$ 0.10 0.020 $\pm$ 0.08 0.602 $\pm$ 0.10 0.692 $\pm$ 0.05 0.373 $\pm$ 0.09 0.693 $\pm$ 0.05
BERT 0.475 $\pm$ 0.12 0.031 $\pm$ 0.09 0.474 $\pm$ 0.12 0.619 $\pm$ 0.08 0.349 $\pm$ 0.09 0.626 $\pm$ 0.08
GPT-2 0.467 $\pm$ 0.17 0.017 $\pm$ 0.08 0.468 $\pm$ 0.17 0.618 $\pm$ 0.08 0.292 $\pm$ 0.10 0.619 $\pm$ 0.08
LM LIME SHAP IG Grad LRP
Distilbert 0.691 $\pm$ 0.05 0.619 $\pm$ 0.06 -0.285 $\pm$ 0.12 -0.215 $\pm$ 0.12 0.706 $\pm$ 0.05
BERT 0.616 $\pm$ 0.08 0.554 $\pm$ 0.09 -0.125 $\pm$ 0.14 -0.123 $\pm$ 0.14 0.639 $\pm$ 0.08
GPT2 0.213 $\pm$ 0.13 0.560 $\pm$ 0.09 0.033 $\pm$ 0.13 0.031 $\pm$ 0.13 0.615 $\pm$ 0.08
Table 10: Correlation with linear Naive Bayes Scores. The scores obtained with SLALOM-eff (value, lin.) are better than those from LIME, SHAP and comparable to LRP scores.
LM values $v$ importances $s$ lin. LIME SHAP LRP
Bert 0.786 $\pm$ 0.01 0.807 $\pm$ 0.01 0.801 $\pm$ 0.01 0.805 $\pm$ 0.01 0.800 $\pm$ 0.01 0.813 $\pm$ 0.01
Distil-BERT 0.688 $\pm$ 0.01 0.681 $\pm$ 0.01 0.686 $\pm$ 0.01 0.702 $\pm$ 0.01 0.668 $\pm$ 0.01 0.703 $\pm$ 0.01
GPT-2 0.674 $\pm$ 0.01 0.685 $\pm$ 0.01 0.683 $\pm$ 0.01 0.632 $\pm$ 0.01 0.671 $\pm$ 0.01 0.699 $\pm$ 0.01
Table 11: Comparison of techniques to predict Human Attention. SLALOM-eff (importances) perform better than SHAP, comparably to LIME and slightly below LRP.
F.4 Insertion and Removal Benchmarks
@@
It is important to verify that SLALOM scores are competitive to other methods in classical explanation benchmarks as well. We therefore ran the classical removal and insertion benchmarks with SLALOM compared to baselines such as LIME, SHAP, Grad (Simonyan et al., 2013), and Integrated Gradients (IG, Sundararajan et al., 2017). For the insertion benchmarks, the tokens with the highest attributions are inserted to quickly obtain a high prediction score to the target class. For the deletion benchmark, the tokens with the highest attributions are deleted from the sample to obtain a low score for the target class. We subsequently delete/insert more tokens and compute the “Area Over the Perturbation Curve” (AOPC) as in DeYoung et al. (2020), which should be high for deletion and low for insertion. In addition to the insertion results in Table 2(c), the removal results are shown in Table 12(a). We show results for the Yelp dataset in Table 12(b) and Table 12(c). We claim that linear SLALOM scores perform on par with LIME and SHAP in this benchmark, but do not always outperform them in this metric. For surrogate techniques (LIME, SHAP, SLALOM) we use 5000 samples each.
SLALOM-fidel SLALOM-eff
LM $v$-scores lin. $v$-scores lin. LIME SHAP IG Grad LRP
BERT 0.893 $\pm$ 0.012 0.901 $\pm$ 0.012 0.881 $\pm$ 0.010 0.885 $\pm$ 0.010 0.875 $\pm$ 0.012 0.881 $\pm$ 0.011 0.084 $\pm$ 0.010 0.069 $\pm$ 0.008 0.852 $\pm$ 0.019
DistilBERT 0.841 $\pm$ 0.014 0.854 $\pm$ 0.013 0.888 $\pm$ 0.008 0.886 $\pm$ 0.008 0.838 $\pm$ 0.012 0.864 $\pm$ 0.009 0.143 $\pm$ 0.012 0.131 $\pm$ 0.012 0.865 $\pm$ 0.011
GPT-2 0.837 $\pm$ 0.013 0.844 $\pm$ 0.013 0.782 $\pm$ 0.013 0.784 $\pm$ 0.012 0.479 $\pm$ 0.024 0.859 $\pm$ 0.012 0.289 $\pm$ 0.021 0.269 $\pm$ 0.020 0.833 $\pm$ 0.013
average 0.857 $\pm$ 0.013 0.866 $\pm$ 0.013 0.851 $\pm$ 0.010 0.852 $\pm$ 0.010 0.731 $\pm$ 0.016 0.868 $\pm$ 0.011 0.172 $\pm$ 0.014 0.156 $\pm$ 0.013 0.850 $\pm$ 0.014
@@
(a) IMDB: Area-Over Perturbation Curve (deletion, higher is better)
SLALOM-fidel SLALOM-eff
LM $v$-scores lin. $v$-scores lin. LIME SHAP IG Grad LRP
BERT 0.015 $\pm$ 0.005 0.011 $\pm$ 0.005 0.011 $\pm$ 0.005 0.011 $\pm$ 0.005 0.017 $\pm$ 0.009 0.012 $\pm$ 0.005 0.224 $\pm$ 0.024 0.214 $\pm$ 0.022 0.010 $\pm$ 0.005
DistilBERT 0.018 $\pm$ 0.006 0.019 $\pm$ 0.006 0.032 $\pm$ 0.009 0.032 $\pm$ 0.009 0.014 $\pm$ 0.005 0.020 $\pm$ 0.005 0.250 $\pm$ 0.027 0.249 $\pm$ 0.026 0.017 $\pm$ 0.009
GPT-2 0.033 $\pm$ 0.007 0.032 $\pm$ 0.007 0.045 $\pm$ 0.007 0.045 $\pm$ 0.007 0.129 $\pm$ 0.019 0.021 $\pm$ 0.004 0.251 $\pm$ 0.024 0.244 $\pm$ 0.024 0.039 $\pm$ 0.007
average 0.022 $\pm$ 0.006 0.021 $\pm$ 0.006 0.029 $\pm$ 0.007 0.029 $\pm$ 0.007 0.053 $\pm$ 0.011 0.018 $\pm$ 0.005 0.242 $\pm$ 0.025 0.236 $\pm$ 0.024 0.022 $\pm$ 0.007
(b) Yelp: Area-Over Perturbation Curve (insertion, lower is better)
SLALOM-fidel SLALOM-eff
LM $v$-scores lin. $v$-scores lin. LIME SHAP IG Grad LRP
GPT-2 0.747 $\pm$ 0.024 0.753 $\pm$ 0.024 0.726 $\pm$ 0.021 0.727 $\pm$ 0.021 0.444 $\pm$ 0.028 0.849 $\pm$ 0.015 0.292 $\pm$ 0.026 0.290 $\pm$ 0.026 0.740 $\pm$ 0.025
BERT 0.657 $\pm$ 0.038 0.667 $\pm$ 0.038 0.865 $\pm$ 0.012 0.863 $\pm$ 0.013 0.797 $\pm$ 0.022 0.859 $\pm$ 0.013 0.249 $\pm$ 0.028 0.281 $\pm$ 0.029 0.855 $\pm$ 0.017
DistilBERT 0.645 $\pm$ 0.033 0.642 $\pm$ 0.033 0.813 $\pm$ 0.017 0.813 $\pm$ 0.018 0.746 $\pm$ 0.025 0.854 $\pm$ 0.013 0.201 $\pm$ 0.026 0.243 $\pm$ 0.028 0.768 $\pm$ 0.024
average 0.683 $\pm$ 0.032 0.687 $\pm$ 0.032 0.801 $\pm$ 0.017 0.801 $\pm$ 0.017 0.663 $\pm$ 0.025 0.854 $\pm$ 0.014 0.247 $\pm$ 0.027 0.271 $\pm$ 0.028 0.788 $\pm$ 0.022
(c) Yelp: Area-Over Perturbation Curve (deletion, higher is better)
Table 12: Additional results for removal/insertion tests: We show results on the IMDB dataset for removal as well as insertion and removal on the Yelp datset.
F.5 Error Analysis for non-transformer models
@@
We also investigate the behavior of SLALOM for models that do not precisely follow the architecture described in the Analysis section of this paper. In the present work, we consider an attribution method that is specifically catered towards the transformer architecture, which is the most prevalent in sequence classification. We advise caution when using our model when the type of underlying LM is unknown. In this case, model-agnostic interpretability methods may be preferred.
However, we investigate this issue further: We applied our SLALOM-eff approach to a simple, non-transformer sequence classification model on the IMDB dataset, which is a three-layer feed-forward network based on a TF-IDF representation of the inputs. We compute the insertion and deletion Area-over-perturbation-curve metrics that are given in Table 13.
SHAP LIME lin. SLALOM-eff
@@
These results show that due to its general expressivity, the SLALOM model also succeeds to provide explanations for non-transformer models that outperform LIME and SHAP in the removal and insertion tests. We also invite the reader to confer Table 2(b), where we show that SLALOM can predict human attention for large models, including the non-transformer Mamba model (Gu & Dao, 2023).
F.6 Runtime analysis